# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

### v2

@@ D.1 BCE-SFT (Reference Model)
The BCE-SFT baseline represents standard supervised fine-tuning.
  • Objective: The model is trained by minimizing the sum of independent Binary Cross-Entropy (BCE) losses for each of the $T$ labels. For an instance $(x_{i},y_{i})$, the loss is $\mathcal{L}_{\text{BCE-SFT}}=\sum_{t=1}^{T}\text{BCE}(m(x_{i};\mathbf{w}_{t}),y_{it})$, where $m(x_{i};\mathbf{w}_{t})$ is the model’s predicted probability for label $t$.
  • Implementation: A separate non-linear classifier head (same MLP as used in FairPO’s main experiments) is trained on top of the frozen ViT features for each label $t$, with parameters $\mathbf{w}_{t}$.
  • Hyperparameters: The primary hyperparameter tuned was the learning rate for the classifier heads. This was selected from $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$ based on the best overall mean Average Precision (mAP) on the validation set. Weight decay was set consistent with FairPO experiments (e.g., 0.01).
  • Role: This model serves two purposes:
    1. 1.
      It acts as a direct performance baseline.
    2. 2.
      The final trained parameters $\{\mathbf{w}_{t}\}$ from this SFT process become the reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}=\{\hat{\mathbf{w}}_{t}\}$ used by the FairPO framework (for DPO-based losses and the non-privileged constraint).
  • Training: Trained for the same number of epochs as FairPO, with early stopping based on overall validation mAP.
@@ D.2 BCE-SFT + Privileged Re-Weighting (RW)
This baseline modifies the standard BCE-SFT by applying static loss re-weighting.
  • Objective: The goal is to improve performance on the privileged labels $\mathcal{P}$ by assigning a higher weight to their contribution in the total loss. The loss function is:
@@
    Typically, $\lambda_{\bar{\mathcal{P}}}$ is set to $1$, and $\lambda_{\mathcal{P}}>1$.
  • Implementation: Same architecture and per-label classifiers as BCE-SFT. The loss computation is modified to incorporate the weights.
  • Hyperparameters:
    • Learning rate: Tuned as for BCE-SFT.
    • Privileged group weight $\lambda_{\mathcal{P}}$: This is the key hyperparameter. It was tuned via grid search on the validation set, selecting from values such as $\{2,3,5,8,10\}$. The value was chosen to maximize mAP on the privileged group $\mathcal{P}$ without causing an excessive drop in mAP on the non-privileged group $\bar{\mathcal{P}}$, or based on the best overall mAP. For our reported results, a weight of $\lambda_{\mathcal{P}}=5$ was found to be effective. $\lambda_{\bar{\mathcal{P}}}$ was fixed at $1$.
  • Training: Trained similarly to BCE-SFT.
@@ D.3 Group DRO + BCE
This baseline applies the Group Distributionally Robust Optimization technique to the standard BCE loss.
  • Objective: To minimize the worst-case BCE loss across the predefined groups (privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$). The objective is:
@@
    where $\mathcal{L}_{\text{BCE}}(\mathcal{G})=\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}\text{BCE}(m(x_{i};\mathbf{w}_{g}),y_{ig})$ is the average BCE loss for labels within group $\mathcal{G}$.
  • Implementation:
    • Architecture: Same as BCE-SFT.
    • Loss Calculation: Per-label BCE losses are computed. These are then aggregated (averaged) separately for all labels in $\mathcal{P}$ to get $\mathcal{L}_{\text{BCE}}(\mathcal{P})$ for the current batch/sample, and similarly for $\bar{\mathcal{P}}$ to get $\mathcal{L}_{\text{BCE}}(\bar{\mathcal{P}})$.
    • Optimization: Model parameters $\{\mathbf{w}_{t}\}$ are updated using AdamW. The group weights $\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}$ are updated iteratively using a mirror ascent step (exponential weighting based on the current group losses, followed by normalization):
      $\alpha_{\mathcal{G}}^{(s+1)}\propto\alpha_{\mathcal{G}}^{(s)}\exp(\eta_{\alpha}\cdot\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{G}))$
      The final loss for backpropagation through model parameters is $\alpha_{\mathcal{P}}^{(s+1)}\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{P})+\alpha_{\bar{\mathcal{P}}}^{(s+1)}\mathcal{L}_{\text{BCE}}^{(s)}(\bar{\mathcal{P}})$.
  • Hyperparameters:
    • Learning rate for model parameters $\{\mathbf{w}_{t}\}$: Tuned as for BCE-SFT.
    • Learning rate for group weights $\eta_{\alpha}$: Tuned from $\{0.01,0.05,0.1,0.2\}$ on the validation set, similar to FairPO’s $\eta_{\alpha}$. The selection aimed for stable convergence and good worst-group performance. For Group DRO with BCE, the raw group losses might be directly used in the exponential update, or a similar loss scaling/normalization approach as discussed for FairPO’s GRPO step could be employed for improved stability if raw loss magnitudes vary significantly between groups or over time. For the reported results, $\eta_{\alpha}$ was set to $0.05$ after tuning.
  • Training: Trained similarly to BCE-SFT and FairPO regarding epochs and batching.
This section provides further details on the experimental setup for the main FairPO variants (FairPO-DPO, FairPO-SimPO, and FairPO-CPO), complementing the information in Section 5 of the main paper.
@@ E.1 Common Setup for All FairPO Variants
Unless specified otherwise, the following setup details apply to all FairPO variants:
  • Base Model: A Vision Transformer (ViT) model, specifically vit-base-patch16-224 [Dosovitskiy et al., 2021], pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k, was used as the feature extractor. The ViT backbone (i.e., the ‘model.vit‘ component) was kept frozen during fine-tuning with FairPO, except for its final encoder layer, which was made trainable to allow for some adaptation of higher-level features.
  • Datasets and Preprocessing: Experiments were conducted on MS-COCO 2014 [Lin et al., 2015] and NUS-WIDE [Chua et al., 2009]. Images were resized to $224\times 224$ pixels and normalized using ImageNet statistics. Standard data augmentations (random horizontal flips, random resized crops) were applied during training. Details on dataset splits and privileged/non-privileged partitioning (20% least frequent labels as privileged) are provided in Section 5 and Appendix C.
  • Optimizer: The AdamW optimizer [Loshchilov and Hutter, 2019] was used for updating the trainable parameters (classifier heads and the last ViT encoder layer).
  • Training Schedule: Models were trained for a maximum of 25 epochs with a batch size of 32. Early stopping was employed based on the overall mAP (average of mAP on privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets) on the validation set, with a patience of 5 epochs.
  • Reference Model ($\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$): The reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$ required by FairPO-DPO and for the non-privileged loss constraint (Eq. 17) in all variants were obtained from a BCE-SFT model (see Appendix D). This SFT model used the same ViT architecture and per-label MLP classifier heads (detailed below) and was trained with a standard BCE loss.
@@ E.2 Per-Label Non-Linear MLP Classifier Head
For each of the $T$ labels in the dataset, a dedicated, independent non-linear Multi-Layer Perceptron (MLP) head was used to predict the probability of that label being positive. This allows each label to have its own specialized classifier operating on the shared image features extracted by the ViT.
  • Architecture: Each MLP head takes the $d$-dimensional output feature vector from the ViT’s [CLS] token (where $d$ is the hidden size of ViT, e.g., 768 for ViT-Base) as input and outputs a single logit for the corresponding label. The architecture of each MLP head is as follows:
@@
      Linear Layer: $d\rightarrow 256$ neurons
@@
      ReLU Activation
@@
      Linear Layer: $256\rightarrow 64$ neurons
@@
      ReLU Activation
@@
      Linear Layer: $64\rightarrow 16$ neurons
    6. 6.
      ReLU Activation
    7. 7.
      Linear Layer: $16\rightarrow 4$ neurons
    8. 8.
      ReLU Activation
    9. 9.
@@
    The final probability score $m(x_{i};\mathbf{w}_{t})$ for label $t$ is obtained by applying a sigmoid function to this output logit.
  • Parameters: The parameters $\mathbf{w}_{t}$ for each label $t$’s MLP head are specific to that label and are learned during the fine-tuning process. The set of all such parameters is denoted by $\{\mathbf{w}_{t}|\forall t\in\mathcal{T}\}$.
  • Trainability: All parameters of these MLP heads were trainable during both SFT (for the reference model) and FairPO fine-tuning.
  • Motivation: Using separate MLP heads rather than a single shared linear layer allows for more complex, non-linear decision boundaries tailored to each label’s specific characteristics and inter-dependencies with the input features. This can be particularly beneficial for labels with varying difficulty or those requiring nuanced feature interpretations.
@@ E.3 FairPO Variant-Specific Hyperparameters
The core FairPO learning rates ($\eta_{\mathbf{w}}$ for model parameters, $\eta_{\alpha}$ for GRPO weights) were tuned via grid search. For $\eta_{\mathbf{w}}$ (AdamW LR for classifier heads and last ViT layer), values were explored from $\{1e-5,5e-5,1e-4,5e-4\}$. For $\eta_{\alpha}$, values from $\{0.01,0.05,0.1,0.2\}$ were explored. The final reported results used $\eta_{\mathbf{w}}=1e-4$ and $\eta_{\alpha}=0.05$ after tuning.
Specific hyperparameters for each FairPO variant’s privileged loss component were tuned as per Table 5 in Appendix F. All hyperparameter tuning was performed on the designated validation set for each dataset. The GRPO mechanism with scaled loss updates (as described in Section 4) was used for adaptively balancing $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$.
@@ F.1 Hyperparameters Tuning
For all experiments, we used the AdamW optimizer [Loshchilov and Hutter, 2019]. The initial learning rate for model parameters ($\eta_{\mathbf{w}}$) was selected from $\{1e-5,5e-5,1e-4,5e-4\}$ and the learning rate for GRPO’s alpha weights ($\eta_{\alpha}$) was selected from $\{0.01,0.05,0.1,0.2\}$. A batch size of 32 was used. The ViT model used was vit-base-patch16-224555https://huggingface.co/docs/transformers/en/model_doc/vit pretrained on ImageNet-21k and fine-tuned on ImageNet-1k. All models were trained for a maximum of 25 epochs with early stopping based on the validation set’s overall mAP (average of $\mathcal{P}$ and $\bar{\mathcal{P}}$ mAP), with a patience of 5 epochs  [^6] All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM)..
Specific hyperparameters for FairPO variants were tuned via grid search on the validation set:
  • FairPO-DPO: $\beta$ (DPO strength parameter, see Eq. 8, 9) was chosen from $\{0.1,0.5,1.0\}$. The non-privileged slack $\epsilon$ (Eq. 17) was chosen from $\{0.01,0.05,0.1\}$.
  • FairPO-SimPO: $\beta$ (SimPO preference term scaling, see Eq. 12) was chosen from $\{0.1,0.5,1.0\}$. The margin $\gamma$ (SimPO, Eq. 12) was chosen from $\{0.05,0.1,0.2\}$. The non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$.
  • FairPO-CPO: $\beta$ (CPO preference term scaling, see Eq. 13) was chosen from $\{0.1,0.5,1.0\}$. The NLL regularizer weight $\lambda_{\text{CPO}}$ (Eq. 14) was chosen from $\{0.1,0.5,1.0\}$. The non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$.
The final selected hyperparameters for each dataset and FairPO variant are reported in Table 5.
Table 5: Final selected hyperparameters for FairPO variants on MS-COCO and NUS-WIDE. Note: For FairPO-CPO, the first value under $\beta/\lambda_{\text{CPO}}$ is $\beta$ and the second is $\lambda_{\text{CPO}}$. Base learning rate $\eta_{\mathbf{w}}$ was $41e-$ 4 and $\eta_{\alpha}$ was $0.05$ for all reported results after tuning.
@@
For the baseline BCE-SFT + Privileged Re-Weighting, the loss weight for privileged labels was set to 5 after tuning from $\{2,3,5,10\}$. For GDRO + BCE, the $\eta_{\alpha}$ was tuned similarly to FairPO.
F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
Table 6 shows the mAP for privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets as $\beta$ varies, keeping other hyperparameters (including $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$) fixed to their optimal values. Performance is relatively stable across a range of $\beta$, though very small or very large values can lead to suboptimal results. Table 7 shows the mAP as $\epsilon$ varies, keeping other hyperparameters (including $\beta=0.5,\lambda_{\text{CPO}}=0.5$) fixed. A moderate $\epsilon$ helps balance the objectives; too small an $\epsilon$ can be overly restrictive on $\bar{\mathcal{P}}$ performance, while too large an $\epsilon$ might allow too much degradation.

### v3

@@ D.1 BCE-SFT (Reference Model)
The BCE-SFT baseline represents standard supervised fine-tuning. Its objective is to minimize the sum of independent Binary Cross-Entropy (BCE) losses for each of the $T$ labels, where the loss for an instance $(x_{i},y_{i})$ is $\mathcal{L}_{\text{BCE-SFT}}=\sum_{t=1}^{T}\text{BCE}(m(x_{i};\mathbf{w}_{t}),y_{it})$. For implementation, a separate non-linear MLP classifier head is trained for each label $t$ on top of frozen ViT features, consistent with the main FairPO architecture. The primary hyperparameter tuned was the learning rate for these classifier heads, selected from $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$ to achieve the best overall mean Average Precision (mAP) on the validation set. Weight decay was set to 0.01, matching the FairPO experiments.
This model serves two critical roles in our study. First, it acts as a direct performance baseline against which we measure fairness improvements. Second, the final trained parameters $\{\mathbf{w}_{t}\}$ from this SFT process become the reference parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}=\{\hat{\mathbf{w}}_{t}\}$ used throughout the FairPO framework, both for the DPO-based losses and for the non-privileged loss constraint. The model was trained for the same number of epochs as FairPO, with early stopping based on overall validation mAP.
@@ D.2 BCE-SFT + Privileged Re-Weighting (RW)
This baseline modifies the standard BCE-SFT by applying static loss re-weighting to improve performance on the privileged labels $\mathcal{P}$. The objective is to assign a higher weight to their contribution in the total loss function:
@@
For this setup, we use the same per-label classifier architecture as BCE-SFT, but the loss computation is modified to incorporate the weights. The key hyperparameter is the privileged group weight $\lambda_{\mathcal{P}}$, which was tuned via grid search from values $\{2,3,5,8,10\}$. The value was chosen to maximize mAP on the privileged group without an excessive drop in performance on the non-privileged group. For our reported results, a weight of $\lambda_{\mathcal{P}}=5$ was found to be effective, while $\lambda_{\bar{\mathcal{P}}}$ was fixed at $1$. Training was conducted similarly to BCE-SFT.
@@ D.3 Group DRO + BCE
This baseline applies the Group Distributionally Robust Optimization technique to the standard BCE loss. The objective is to minimize the worst-case BCE loss across the predefined privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) groups:
@@
Here, $\mathcal{L}_{\text{BCE}}(\mathcal{G})$ is the average BCE loss for all labels within a group $\mathcal{G}$. The implementation uses the same BCE-SFT architecture. Per-label BCE losses are computed and then aggregated separately for the $\mathcal{P}$ and $\bar{\mathcal{P}}$ groups. Model parameters $\{\mathbf{w}_{t}\}$ are updated with AdamW, while the group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are updated iteratively using a mirror ascent step: $\alpha_{\mathcal{G}}^{(s+1)}\propto\alpha_{\mathcal{G}}^{(s)}\exp(\eta_{\alpha}\cdot\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{G}))$. The final loss for backpropagation is the adaptively weighted sum of the group losses.
For hyperparameters, the model learning rate was tuned as for BCE-SFT. The group weight learning rate $\eta_{\alpha}$ was tuned from $\{0.01,0.05,0.1,0.2\}$, with $\eta_{\alpha}=0.05$ selected for its stable convergence and good worst-group performance. For the reported results, raw group losses were used in the exponential update, though we note that a loss scaling approach similar to that in FairPO could also be employed for enhanced stability. The training protocol regarding epochs and batching was identical to FairPO.
@@ E.1 Common Setup for All FairPO Variants
Unless specified otherwise, a common setup was used for all FairPO variants to ensure fair comparison. The base model for feature extraction was a Vision Transformer (ViT), specifically vit-base-patch16-224 [Dosovitskiy et al., 2021], which was pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k. During our fine-tuning, the ViT backbone was kept frozen, with the exception of its final encoder layer, which was made trainable to allow for adaptation of higher-level features. All experiments were conducted on the MS-COCO 2014 [Lin et al., 2015] and NUS-WIDE [Chua et al., 2009] datasets. Images were resized to $224\times 224$ pixels, normalized using ImageNet statistics, and augmented with standard techniques like random horizontal flips and resized crops. The AdamW optimizer [Loshchilov and Hutter, 2019] was used to update all trainable parameters. Models were trained for a maximum of 25 epochs with a batch size of 32, and we employed an early stopping strategy with a patience of 5 epochs based on the overall mAP on the validation set. The reference model parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$, required for FairPO-DPO and the non-privileged loss constraint in all variants, were obtained from a BCE-SFT model detailed in Appendix D.
@@ E.2 Per-Label Non-Linear MLP Classifier Head
For each of the $T$ labels in a dataset, we employed a dedicated and independent non-linear Multi-Layer Perceptron (MLP) head to predict the probability of that label being positive. Using separate MLP heads allows for more complex, non-linear decision boundaries tailored to each label’s specific characteristics, which is particularly beneficial for labels with varying difficulty. Each MLP head takes the $d$-dimensional feature vector (where $d=768$ for ViT-Base) from the ViT’s [CLS] token as input and outputs a single logit. The final probability score $m(x_{i};\mathbf{w}_{t})$ is obtained by applying a sigmoid function to this logit. The specific architecture for each MLP head is as follows:
@@
    Linear Layer: $d\rightarrow 256$ neurons, followed by ReLU Activation
@@
    Linear Layer: $256\rightarrow 64$ neurons, followed by ReLU Activation
@@
    Linear Layer: $64\rightarrow 16$ neurons, followed by ReLU Activation
@@
    Linear Layer: $16\rightarrow 4$ neurons, followed by ReLU Activation
@@
The parameters $\mathbf{w}_{t}$ for each label $t$’s MLP head are unique to that label. All parameters within these MLP heads were fully trainable during both the SFT pre-training (for the reference model) and the final FairPO fine-tuning.
@@ E.3 FairPO Variant-Specific Hyperparameters
The core FairPO learning rates, $\eta_{\mathbf{w}}$ for model parameters and $\eta_{\alpha}$ for GRPO weights, were tuned via grid search. For $\eta_{\mathbf{w}}$ (the AdamW LR), values were explored from $\{1e-5,5e-5,1e-4,5e-4\}$, and for $\eta_{\alpha}$, values were explored from $\{0.01,0.05,0.1,0.2\}$. After tuning, the final reported results used $\eta_{\mathbf{w}}=1e-4$ and $\eta_{\alpha}=0.05$. The specific hyperparameters for each FairPO variant’s privileged loss component (e.g., $\beta$, $\gamma$, $\lambda_{\text{CPO}}$) were also tuned on the designated validation set for each dataset, with final values reported in Table 5 (Appendix F). Throughout all experiments, the GRPO mechanism with scaled loss updates, as described in Section 3, was used to adaptively balance the privileged and non-privileged loss terms.
@@ F.1 Hyperparameters Tuning
For all experiments, we used the AdamW optimizer [Loshchilov and Hutter, 2019] with a batch size of 32. The base model was a vit-base-patch16-224222https://huggingface.co/docs/transformers/en/model_doc/vit pretrained on ImageNet-21k and fine-tuned on ImageNet-1k. The initial learning rate for model parameters ($\eta_{\mathbf{w}}$) was selected from $\{1e-5,5e-5,1e-4,5e-4\}$, while the learning rate for GRPO’s alpha weights ($\eta_{\alpha}$) was selected from $\{0.01,0.05,0.1,0.2\}$. All models were trained for a maximum of 25 epochs with early stopping based on the validation set’s overall mAP, using a patience of 5 epochs [^3] All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM)..
Specific hyperparameters for each FairPO variant were tuned via grid search on the validation set. For FairPO-DPO, the strength parameter $\beta$ was chosen from $\{0.1,0.5,1.0\}$, and the non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$. For FairPO-SimPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the margin $\gamma$ from $\{0.05,0.1,0.2\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$. For FairPO-CPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the NLL regularizer weight $\lambda_{\text{CPO}}$ from $\{0.1,0.5,1.0\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$. The final selected hyperparameters for each dataset and FairPO variant are reported in Table 5.
For our baselines, the loss weight for privileged labels in BCE-SFT + Privileged Re-Weighting was set to 5 after tuning from $\{2,3,5,10\}$. The group weight learning rate $\eta_{\alpha}$ for GDRO + BCE was tuned similarly to FairPO.
F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
Table 5: Final selected hyperparameters for FairPO variants on MS-COCO and NUS-WIDE. Note: For FairPO-CPO, the first value under $\beta/\lambda_{\text{CPO}}$ is $\beta$ and the second is $\lambda_{\text{CPO}}$. Base learning rate $\eta_{\mathbf{w}}$ was 1e41e-4 and $\eta_{\alpha}$ was $0.05$ for all reported results after tuning.
@@ F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
Table 6 shows the mAP for privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets as $\beta$ varies, keeping other hyperparameters (including $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$) fixed to their optimal values. Performance is relatively stable across a range of $\beta$, though very small or very large values can lead to suboptimal results.
@@
Table 7 shows the mAP as $\epsilon$ varies, keeping other hyperparameters (including $\beta=0.5,\lambda_{\text{CPO}}=0.5$) fixed. A moderate $\epsilon$ helps balance the objectives; too small an $\epsilon$ can be overly restrictive on $\bar{\mathcal{P}}$ performance, while too large an $\epsilon$ might allow too much degradation.